OpenStudy (anonymous):

A bus route shows a very regular pattern one morning: 8 passengers get on at the first stop. From then on, 3 passengers get off and 5 passengers get on at every stop, until the entire bus is full. If the bus holds 45 passengers in all, at which stop does the bus fill up?

OpenStudy (mertsj):

1. 8 2. 8-3+5=10 3. 10-3+5=12 4. 12-3+5=14 5. 14-3+5=16 Are we seeing a pattern here? Can you solve the problem now?

OpenStudy (anonymous):

oh okay so its adding 2

OpenStudy (mertsj):

Yes. If 3 people get off and 5 get on, that is a net gain of 2 people at each stop after the first one.

OpenStudy (anonymous):

Oh okay but what would be an equation expressing this situation

OpenStudy (mertsj):

8+2x>=45

OpenStudy (anonymous):

why >= ??

OpenStudy (mertsj):

Because it won't come out exactly on 45 and you don't want to have half of a stop.

OpenStudy (anonymous):

but the teacher said the answer is like the 20th?

OpenStudy (mertsj):

Some people have to do everything the hard way. 1. 8 2. 8-3+5=10 3. 10-3+5=12 4. 12-3+5=14 5. 14-3+5=16 6. 16-3+5=18 7. 18-3+5=20 8. 20-3+5=22 9. 22-3+5=24 10. 24-3+5=26 11. 26-3+5=28 12. 28-3+5=30 13. 30-3+5=32 14. 32-3+5=34 15. 34-3+5=36 16. 36-3+5=38 17. 38-3+5=40 18. 40-3+5=42 19. 42-3+5=44 20. 44-3+5=46 Walla!! On the 20th stop, they get to the 45th passenger. 8+2x>=45 2x>=37 x>=18.5 so 19 Add 1 for the 1st stop. Walla!!! On the 20th stop they get to the 45th passenger.

OpenStudy (anonymous):

wait what do you mean by add 1 to the first stop O:

OpenStudy (mertsj):

Did I say to add 1 to the first stop?

OpenStudy (anonymous):

I ment like what do you mean by add 1 for the first stop

OpenStudy (mertsj):

After the first stop, how many people do they need to get to 45 people?

OpenStudy (anonymous):

36

OpenStudy (mertsj):

How do you get 36?

OpenStudy (anonymous):

45-8?

OpenStudy (anonymous):

just kidding, 37 lol

OpenStudy (mertsj):

How many people do they gain at each stop?

OpenStudy (anonymous):

2